ONE NUMBER IS 8 MORE THAN A SECOND NUMBER. THREE TIMES THE SMALER NUMBER PLUS TWICE THE LARGER IS EQUAL TO 26. WHAT IS THE TWO NUMBER=
Answer by lwsshak3(11628)

ONE NUMBER IS 8 MORE THAN A SECOND NUMBER. THREE TIMES THE SMALLER NUMBER PLUS TWICE THE LARGER IS EQUAL TO 26. WHAT IS THE TWO NUMBER
..
let x=smaller number
x+8 = larger number
3x+2(x+8)=26
3x+2x+16=26
5x=26-16=10
x=10/5=2
x+8=10
ans:
smaller number=2
larger number=10